x-4 a) lim x-2 X-2 xe* b) lim *-01-et lim (1+x)* X+00
Answer #1


) (e 샤 르 - 24 212 - 가 29 - <a la) (a)) 가 222 (29) 2-2 - Got) 92 = 2+ ed 미르 = 4 292 2-2 e It ne? ato The (응 Sorm) by ho Hospit